# CARTE: pretraining and transfer for tabular learning

### v1

@@ 3.3 Fine-tuning for downstream tasks
From the architecture shown in Figure 3, the fine-tuning strategy for CARTE is to take two parts from the pretrained model: the initial layers for nodes and edges (blue and red blocks) and the ‘Aggregation & Readout’ layer. Although such simplification is contrary to conventional fine-tuning approaches, it is a suitable one in the view of graph-neural networks. Recall that the graph of downstream table entities would be star-like graphs (Figure 1). Given that the downstream tables contain less variability in graph structures and the cardinality of discrete variables compared to YAGO, the latent representations after each attention layer would lose discriminant characteristics (also known as the over-smoothing problem, Chen et al., 2020a; Rusch et al., 2023, which we study in appendix C.1). Therefore, we take the widely used convention of setting the number of attention layers as the maximum $k$-hop relation, which corresponds to $k=1$. For the final classifier, we simply attach the linear layers. With the base model for fine-tuning, we consider two different settings of downstream inference.
@@
This is the well-known setting in which we are given a single table with a target variable to predict. Before transforming table entities into graphs, we preprocess numerical variables with a power transform (Yeo & Johnson, 2000). The power transform has shown to be effective in several works (e.g. Hollmann et al., 2023; Cvetkov-Iliev et al., 2022), and likewise, gives stability to the fine-tuning process of CARTE. Moreover we employ a bagging strategy (Breiman, 1996), in which different models, based on different train-validation splits used for early-stopping, are trained. The prediction outputs are calculated as the average of the outputs from each model.
@@
We also use CARTE in transfer learning setting where we are given a source table $X_{S}$ that can aid predictions on our target table $X_{T}$. Importantly, the source table may have many more samples than the target. We fine-tune CARTE on both tables jointly. Note that the graph representation of the tables enables such joint fine tuning without correspondences in the columns; however, we do need to have similar outcomes $y_{S}$ and $y_{T}$ on both tables. The source outcomes $y_{S}$ are transformed to match the first moment of the target outcome $y_{T}$ using as above a power transform (Yeo & Johnson, 2000, note that here we use the inverse transform). If the target and source table differ on the classification / regression nature of the outcome, we adapt $y_{S}$ as the following: for a classification target $y_{T}$ we binarize regression outcomes in the source table, and for the target regression $y_{T}$, we use binary classification outcomes of the source table, encoded as $\{0,1\}$ and standard scaled. We then proceed to fine tune CARTE by drawing a batch with a fixed proportion of rows from the target and the source table (we use a batch size of 64, 8 of which come from the target). As we consider the settings in which the target table has much fewer samples than the source table, we iterate many more times on the target table, but do early stopping on a validation set of the target table. Note that as in the single-table case, we use a bagging strategy of building multiple learners on different random validation sets and averaging the predictions. Often, early stopping kicks in before all the data points of the source have been covered. This prevents overfitting the source data, which may be less important than the target data for predicting $y_{T}$. Moreover, we use the hyperparameters that have been selected above for the single-table setting.
As we choose source tables quite loosely from weakly-related data, the resulting pair-wise learner may not actually improve upon the single-table learner if the source table does not bring in enough related information. We thus combine the pair-wise learner with the single-table learner ensembling the predictors by combining their output with a softmax. The weights of the softmax are computed using the prediction score computed in the internal validation set of these learners, but divided by the standard deviation across the learners to set the temperature of the softmax.
@@
The key to transfer learning, as above, is finding the right source table. If we have multiple tables from a given domain or institution, CARTE can be adapted to use them all, finding the most useful information for transfer. In these settings we are given a target table $X_{T}$ and a set of source tables ${X_{S,1}...X_{S,m}}$. We proceed to build individual learners: first the single-table learner on $X_{T}$, then each pair [^2] To limit computation cost, we do not explore the full combinatorials of source tables. of one source table $X_{S,i}$ and the target table $X_{T}$ using the pair-wise joint learner described above. Here again, not every pair-wise learning brings the same amount of useful information. Thus, to find the optimal combination of datasets, we use the same strategy as above of ensembling all the pair-wise predictors as well as the single-table predictor. As a consequence, if all source tables lead to predictors that work as well, they are combined with equal weights, but if one source dominates, the prediction is anchored on this one.
@@ 4.1 Experimental setup
We used 51 tabular learning datasets, all with an associated learning task –40 regressions and 11 classification–, gathered across multiple sources, mainly from previous machine learning studies and kaggle competitions. These datasets cover a variety of topics of society and businesses: accidents, elections, remunerations, food, restaurants, etc. We select datasets that were representative of modern data science applications: tabular data with meaningful columns and discrete entries, unlike many datasets from UCI . Appendix B gives the specific list of datasets.
@@
We evaluate different baselines, with the following abbreviations (details and hyper-parameter tuning in Appendix A.2):
TabVec-…-HGB,
a family of baselines using the TableVectorizer from the Skrub package (Skrub, 2024) to encode tables into numerical arrays. When applied on all features (TabVec-HGB), text features are encoded using the Gamma-Poisson encoder from skrub, introduced in Cerda & Varoquaux (2022), which extracts latent categories from substrings. We also use tailored strategy to encode columns that we manually identify as containing rich strings or entity names. For this the corresponding columns are encoded either with FastText (Joulin et al., 2017) (TabVec-FT-HGB) or with an LLM –we use intfloat/e5-small-v2 (Wang et al., 2022) from HuggingFace– (TabVec-FT-LLM, and the TableVectorizer is applied on remaining numerical and categorical features. For all these baselines, the encoded table is passed into the HistGradientBoosting estimators from scikit-learn (Pedregosa et al., 2011).
@@
(Dorogush et al., 2018), a gradient-boosted trees package commonly used to learn on tables. We treat text features as categorical, which are encoded by CatBoost’s categorical encoding, an improved version of target encoding (Micci-Barreca, 2001).
@@
Inspired by TabLLM (Hegselmann et al., 2023), we investigate encoding each row of a table with an LLM. We represent each row as a sentence, alternate the column names and the corresponding entries, and encode the table with the last hidden layer. Unlike TabLLM, however, the encoded table is passed to the HistGradientBoosting estimator as this led to better prediction accuracies. The benefit of using an LLM is that, as with CARTE, it gives a consistent representation of data without requiring correspondences in columns. For numerical entries, we investigate either concatenating them as additional features outside of the LLM (CN), or passing them as strings to the LLM (EN).
MLP
The classical Multilayer Perceptron with target encoder (Micci-Barreca, 2001) for treating categorical features. The missing values were imputed with the mean for numerical features, while they were treated as another category for categorical features. After the transformations, we applied the minmax scale on all columns to set the values in between zero and one.
Resnet
Similar architecture to the MLP (with the same encoding and transformation measures), but with additional layernorm/batchnorm and skip-connections.
@@
a. Regression – 40 datasets HGB = sklearn’s HistGradientBoosting TabVec = skrub’s TableVectorizer FT = FastText CN = concat numerical EN = embed numerical                                                                 Refer to captionb. Classification – 11 datasets
Refer to captionRefer to caption
Figure 4: CARTE outperforms alternatives for learning on single tables. Learning curve on (a) regression and (b) classification tasks over 51 datasets. Top: normalized score (for a given dataset, 1 is the best performing score across all methods and training-set size, and 0 is the worst), averaged across all datasets. Bottom: critical difference diagrams, for $n=1024$ (code from Ismail Fawaz et al., 2019).
@@ 4.2 Results on single tables
Figure 4 compares the prediction performance of the multiple methods, summarizing the different datasets. We see that CARTE consistently outperforms alternatives across the different sample sizes, whether it is with normalized score [^3] The calculation of the normalized score was adapted from Grinsztajn et al. (2022), in which the minimum score is fixed at $\rho$ ($\rho=0$ for regression and $\rho=0.5$ for classification). or critical difference diagrams based on the Wilcoxon-Holm method to detect pairwise significance.
@@
The hypothesis to explain the good performance of CARTE is that has integrated information on many entities because it has been pretrained on the YAGO knowledge base. However, in a given downstream table, entities might appear written in a different way: for instance ”Londres” instead of ”London”. This begs the question of whether CARTE transfers well useful information if the string representation of the entities differs. String matching is necessary for instance when using as features vector embeddings of the entries such as KEN embeddings (Cvetkov-Iliev et al., 2023).
Figure 5: Entity matching not required for CARTE, and downstream entities do not need in YAGO to work. We evaluate CARTE and KEN either on the full datasets, or on a reduced version of the datasets corresponding to entities present in YAGO. In addition, when entities are present in YAGO, we compare matching them to their canonical name in YAGO (blue) or leaving the original names (orange). When KEN is used enrich the dataset, CatBoost is used for prediction, and entities without match are replaced with missing values. Each point on the figure correspond to an improvement in performance with respect to Catboost without any enrichment for a dataset and a training size. Detailed results in Appendix C.2.
On four of our datasets (company employees, movies, US accidents, and US election), we performed manual entity matching of the entries to their corresponding YAGO entity. Figure 5 shows that while using KEN requires entity matching to have good performance, CARTE has string-level models, and its performance does not vary between using original entries or manually matched entities.
@@ 4.3 Learning across multiple tables
Joint learning84 cases of two tables                                                                                  HGB = sklearn’s HistGradientBoosting UM = columns unmatched M = columns matched
@@
We investigate learning across multiple tables without explicit correspondences across columns. We use tables “in the wild”: in our 51 datasets, we find groups covering the same general topic (selling bikes, rating restaurants) though they come from different sources (appendix B.3). In this setting, we can readily use CARTE and S-LLM approaches as they use an open-vocabulary representation of the column to embed entries (but only the EN, Embed Numerical, version of S-LLM, as the CN, Concat Numerical, needs correspondence for the numerical columns). As CatBoost natively deals with missing values, we use it by concatenating the datasets and adding missing values for columns that mismatch. We also investigate manual matching of columns.
@@
Figure 7: CARTE further benefits from additional source tables. Each point represents an average increase in the performance of CARTE on a target table from the case of no source tables, when incorporating different number of sources into its training process. For each setting, we permute the tables so that each table is used both as target and as source.
@@
The difficulty for transfer learning may be finding good source tables. In Figure 7, we investigate bringing in more source tables, up to a total of 4 tables (1 target, and 3 sources, with a total of 245 cases). CARTE benefits from adding source tables: not only does the median performance improve but also, the variability with respect the lower range improves quickly. In other words, with more source tables, there are more chances of finding a real beneficial one, and thus the worst-case scenario becomes better.
5 Conclusion
By using an architecture that models table entries not as a $i^{\text{th}}$ feature in a data matrix but as a function of its context (column names and neighboring entries) as well open-vocabulary embedding of strings, CARTE enables consistent representations of very different tabular data. This opens the door to pre-training across background tables, and fine-tuning to downstream tasks without need for entity or schema matching. Our results show that after pretraining on a large knowledge base, the resulting model brings marked benefits to downstream analytic tasks, consistently outperforming a broad range of baselines both for learning on a single table or transfer learning on tables with imperfect correspondences. It enables transfer from tables in the wild, a setting so far never studied. Pre-training has been key to the wide application of deep learning on images and text. We hope that the ideas behind CARTE will bring these benefits to tabular learning, enabling further improved architectures, for instance by merging with the complementary meta-learning ideas of TabPFN (Hollmann et al., 2023).

### v2

@@ 3.3 Fine-tuning for Downstream Tasks
For a given downstream task, fine-tuning CARTE proceeds by reusing only part of the pretrained architecture (as shown in Figure 3): the initial layers for nodes and edges (blue and red blocks) and the ‘Aggregation & Readout’ layer. Though such simplification differs from many fine-tuning approaches, it stems from the behavior of graph-neural networks. Indeed, downstream table entities form simpler graphs than during pre-training. First they are star-like (Figure 1). Second, the downstream tables contain less variability in graph structures and less cardinality of discrete variables compared to YAGO. Too deep an architecture risks washing out discriminant characteristics in the output representations (the over-smoothing problem, Chen et al., 2020a; Rusch et al., 2023, studied in Appendix C.4). Therefore, we use a convention common in graph models: setting the number of attention layers as the maximum $k$-hop relation, here $k=1$. For the final classifier, we simply attach the linear layers. With the base model for fine-tuning, we consider two different settings of downstream inference.
@@
This is the well-known setting in which we are given a single table with a target variable to predict. Before transforming table entities into graphs, we preprocess numerical variables with a power transform (Yeo & Johnson, 2000). The power transform has shown to be effective in several works (e.g., Hollmann et al., 2023; Cvetkov-Iliev et al., 2023), and likewise, gives stability to the fine-tuning process of CARTE. Moreover we employ a bagging strategy (Breiman, 1996), in which different models, based on different train-validation splits used for early-stopping, are trained. The prediction outputs are calculated as the average of the outputs from each model.
@@
We also use CARTE in transfer learning settings where we are given a source table $X_{S}$ that can aid predictions on our target table $X_{T}$. Importantly, the source table may have larger train samples than the target. We fine-tune CARTE on both tables jointly. The graph representation enables such joint fine tuning without correspondences in the columns; however, we do need to have similar outcomes $y_{S}$ and $y_{T}$ on both tables. The source outcomes $y_{S}$ are transformed to match the first moment of the target outcome $y_{T}$ using as above a power transform (Yeo & Johnson, 2000, note that here we use the inverse transform). If the target and source table differ on the classification / regression nature of the outcome, we adapt $y_{S}$ as the following: for a classification target $y_{T}$, we binarize regression outcomes in the source table, and for a regression target $y_{T}$, we use binary classification outcomes of the source table, encoded as $\{0,1\}$ and standard scaled. We then proceed to fine tune CARTE by drawing batches with a fixed proportion of rows from the target and source tables (we use a batch size of 64, 8 of which come from the target). We use early stopping on a validation set of the target table, and still rely on the bagging strategy of building multiple learners on different random validation sets and averaging the predictions. Often, early stopping kicks in before all the data points of the source have been covered. This prevents overfitting the source data, which may be less important than the target data for predicting $y_{T}$. We use the hyperparameters selected in the single-table setting.
As we choose source tables quite loosely from weakly-related data, the resulting pairwise learner may not actually improve upon the single-table learner if the source table does not bring in enough related information. We thus combine the pairwise learner with the single-table learner ensembling the predictors by combining their output with a softmax. The weights of the softmax are computed using the prediction score computed in the internal validation set of these learners, but divided by the standard deviation across the learners to set the temperature of the softmax.
@@
The key to transfer learning, as above, is finding the right source table. If we have multiple tables from a given domain or institution, CARTE can be adapted to use them all, finding the most useful information for transfer. In these settings we are given a target table $X_{T}$ and a set of source tables ${X_{S,1}...X_{S,m}}$. We proceed to build individual learners: first the single-table learner on $X_{T}$, then each pair [^2] To limit computation cost, we do not explore the full combinatorials of source tables. of one source table $X_{S,i}$ and the target table $X_{T}$ using the pairwise joint learner described above. Here again, not every pairwise learning brings the same amount of useful information. Thus, to find the optimal combination of datasets, we use the same strategy as above of ensembling all the pairwise predictors as well as the single-table predictor. As a consequence, if all source tables lead to predictors that work as well, they are combined with equal weights, but if one source dominates, the prediction is anchored on this one.
@@ 4.1 Experimental Setup
We use 51 tabular learning datasets, all with an associated learning task –40 regressions and 11 classification–, gathered across multiple sources, mainly from previous machine learning studies and kaggle competitions. They cover a variety of topics of society and businesses: accidents, elections, remunerations, food, restaurants, etc. We select datasets representative of modern data science applications: tables with meaningful columns and discrete entries (Table 3), unlike many datasets from UCI . Appendix B gives the specific list of datasets.
@@
We evaluate different baselines with the following abbreviations (specific details on experiment settings and hyper-parameter tuning are presented in Appendix A.2):
@@
(Dorogush et al., 2018) A gradient-boosted trees package commonly used to learn on tables. We treat text features as categorical, encoded by CatBoost’s categorical encoding, an improved version of target encoding (Micci-Barreca, 2001).
TabVec
The TableVectorizer from the Skrub package (Skrub, 2024) to encode tables that contain string entries into numerical arrays. Columns with low cardinality (number of categories) are one-hot encoded while those with high cardinality are encoded using the Gamma-Poisson encoder from skrub, introduced in Cerda & Varoquaux (2022), which extracts latent categories from substrings. For non tree-based models, missing values are imputed with the mean for numerical features, and treated as another category for categorical features. For neural network models, minmax scale is applied to set values between zero and one.
XGB, HGB, and RF
Tree-based models: XGBoost (Chen & Guestrin, 2016), HistGradientBoosting and RandomForest (from scikit-learn, Pedregosa et al., 2011).
MLP and ResNet
The classical Multilayer Perceptron (MLP) and its extension with additional layernorm/batchnorm and skip-connections (ResNet).
Ridge and Logistic
Linear models, Ridge and Logistic regression for regression and classification tasks.
@@
Inspired by TabLLM (Hegselmann et al., 2023), we investigate encoding each row of a table with a large language model (LLM). We represent each row as a sentence, and encode them with intfloat/e5-small-v2 (Wang et al., 2022) from HuggingFace. Unlike TabLLM, however, the encoded table is passed to the XGB estimator to enable learning for both regression and classification. For numerical entries, we investigate either concatenating them as additional features outside of the LLM (CN), or passing them as strings to the LLM (EN).
@@
Figure 4: CARTE performs best for learning on single tables. Learning curve on (a) regression and (b) classification tasks. Top: normalized score (1 is the best performer across all methods and train size for a dataset, and 0 the worst), averaged across datasets. Bottom: critical difference diagrams (Terpilowski, 2019), for all train sizes. Figure 9 gives critical difference diagram for all methods studied.
@@ 4.2 Results on Single Tables
Figure 4 compares the prediction performance of the multiple methods, summarizing the different datasets. We see that CARTE consistently outperforms alternatives across the different sample sizes, whether it is with normalized score [^3] The calculation of the normalized score was adapted from Grinsztajn et al. (2022), in which the minimum score is fixed at a value $\rho$: $\rho=0$ for regression and $\rho=0.5$ for classification. or critical difference diagrams based on the Conover post hoc test after the Friedman test to detect pairwise significance (Conover, 1999). Another important point is that the bagging strategy used for CARTE also has positive impacts for neural network models: such a bagging with different train/validation splits for early stopping may be beneficial for deep learning in general. Appendix C.1 gives comprehensive results of CARTE and 42 baselines.
CARTE is robust to missing values.
When handling missing values, CARTE discards columns with the missing value in the graph construction step. For example, a data point with one missing value on a table with 10 columns would have nine leaf-nodes after the graph construction step. Table 1 compares the percentage drop in performance and its normalized scores (in comparison to Figure 4) of CARTE and several decision tree baselines that inherently handle the missing values. In the experiment, we randomly drop a proportion of features for each sample (train/test inclusive). The fraction of dropped columns are set as 0.1 (10 % features dropped) and 0.3. The table shows that CARTE continues to outperform the baselines with smaller decrease in performance created by missing values.
Table 1: CARTE is robust to missing values. Percentage drop in performance and its normalized scores with missing values in which a proportion (0.1 or 0.3) of features are randomly dropped.
Percentage decrease created by missing values
Methods Train size (Missing fraction)
64 (0.1) 64 (0.3) 512 (0.1) 512 (0.3)
CARTE 13.28% 38.35% 10.19% 24.42%
CatBoost 21.70% 53.32% 12.23% 29.70%
TabVec-XGB 15.11% 51.27% 12.61% 30.35%
TabVec-RF 7.68% 44.43% 12.77% 29.79%
Normalized absolute score (as in Figure 4)
CARTE 0.44${}_{\textbf{(0.20)}}$ 0.29${}_{\textbf{(0.18)}}$ 0.75${}_{\textbf{(0.12)}}$ 0.61${}_{\textbf{(0.14)}}$
CatBoost 0.31${}_{\text{(0.22)}}$ ${}_{\text{(0.17)}}$ 0.65${}_{\text{(0.15)}}$ 0.50${}_{\text{(0.15)}}$
TabVec-XGB 0.19${}_{\text{(0.20)}}$ 0.11${}_{\text{(0.15)}}$ 0.65${}_{\text{(0.17)}}$ 0.50${}_{\text{(0.17)}}$
TabVec-RF 0.23${}_{\text{(0.21)}}$ 0.14${}_{\text{(0.16)}}$ 0.63${}_{\text{(0.15)}}$ 0.49${}_{\text{(0.15)}}$
Computation time trade-off for CARTE
Table 2 shows the average computation time (in seconds) of the top-four baselines. The strong prediction performance of CARTE comes at a cost in computation time, and the gap increases with train size ($n$). This cost calls for further optimizations.
Table 2: Computation time (in seconds) for top-four baselines for preprocessing, training and testing, across the 51 datasets.
Methods Preprocessing Learning Learning
$n=64$ $n=512$
CARTE 50.20$\pm$63.68 85.43$\pm$60.30 315.49$\pm$119.84
CatBoost       - 0.98$\pm$1.19      1.05$\pm$1.06
TabVec-XGB 64.72$\pm$139.23 0.40$\pm$0.21      1.19$\pm$0.94
S-LLM-XGB 207.87$\pm$361.56 0.87$\pm$0.71      3.49$\pm$1.79
@@
The hypothesis to explain the good performance of CARTE is that has integrated information on many entities because it has been pretrained on the YAGO knowledge base. However, in a given downstream table, entities might appear written in a different way: for instance “Londres” instead of “London”. This begs the question of whether CARTE transfers well useful information if the string representation of the entities differs. String matching is necessary for instance when using as features vector embeddings of the entries such as KEN embeddings (Cvetkov-Iliev et al., 2023).
Figure 5: Entity matching not required for CARTE, and downstream entities do not need to be in YAGO. We evaluate CARTE and KEN either on the full datasets, or on a reduced version of the datasets corresponding to entities present in YAGO. In addition, when entities are present in YAGO, we either match them to their canonical names in YAGO (blue) or keep the original names (orange). When KEN is used to enrich the dataset, CatBoost is used as the estimator, and entities without matching are replaced with missing values. Each point on the figure correspond to an improvement in performance with respect to Catboost without any enrichment. That KEN brings performance gains to CatBoost on YAGO entities confirms the added value of background information. Appendix C.5 gives detailed results.
On four of our datasets (company employees, movies, US accidents, and US election), we performed manual entity matching of the entries to their corresponding YAGO entity. Figure 5 shows that while using KEN requires entity matching to have good performance, the string-level modeling in CARTE make its performance robust to entity variants: using manually matched entities or original entries. Ablations confirm the importance of string-level representations that also capture semantic similarity (Appendix Figure 10).
Comparison to TabLLM baselines
We compare CARTE to three baselines over nine datasets presented TabLLM (Hegselmann et al., 2023). Compared to the datasets in Figure 4, the datasets in TabLLM contain higher fraction of numerical features with less cardinality of categorical columns (see Table 3). Figure 6 gives the critical difference diagram of the baselines. TabPFN shows strength in such settings. Yet, CARTE can attain competitive performances although it is geared towards handling both numerical values and strings. Detailed results are analyzed in Appendix C.2.
Table 3: Difference between this study’s benchmark and TabLLM datasets. Our benchmark datasets contain more categorical columns, in particular with higher cardinality ($|C|$).
Characteristics This study’s TabLLM
benchmark
Fraction of numerical cols. 0.194 0.613
Fraction of cols. with $|C|>10$ 0.625 0.043
Cardinality over data size 0.263 0.001
Figure 6: Comparison to three baselines in TabLLM (Hegselmann et al., 2023). The datasets contain mostly numerical features or low-cardinality categorical columns. In such settings, TabPFN performs best, followed by CARTE, XGBoost, and TabLLM.
@@ 4.3 Learning Across Multiple Tables
We investigate learning across multiple tables without explicit correspondences across columns. We use tables “in the wild”: in our 51 datasets, we find groups covering the same general topic (bike prices, restaurant ratings) though they come from different sources (Appendix B.3). In this setting, we can readily use CARTE and S-LLM approaches as they use an open-vocabulary representation of the column to embed entries (but only the EN, Embed Numerical, version of S-LLM, as the CN, Concat Numerical, needs correspondence for the numerical columns). As CatBoost natively deals with missing values, we use it by concatenating the datasets and adding missing values for mismatch columns. We also investigate manual matching of columns.
@@
The difficulty for transfer learning may be finding good source tables. In Figure 8, we investigate bringing in more source tables, up to a total of 4 tables (1 target, and 3 sources, with a total of 245 cases). CARTE benefits from adding source tables: not only does the median performance improve but also, the lower bound in variability improves. In other words, more source tables give higher chances of finding a beneficial one, and thus the worst-case scenario becomes better.
Figure 8: CARTE further benefits from additional source tables.
5 Discussion and Conclusion
Strings and numbers in tables
Our study touches on the importance of strings in tabular data. They are often overlooked in tabular machine learning (Table 3 shows how datasets are mostly numerical or low-cardinality categories), but central to database research, which focuses on discrete entries. Compared to most tabular-learning models (whether tree-based, or neural networks as TabPFN), the edge of CARTE is on the strings. String preprocessing in skrub’s TableVectorizer also boosts baselines (Figure 9). Conversely, language models, as LLMs, focus on strings, enabling pretraining on huge corpora. They give great preprocessing of strings but must be combined with tree-based methods to handle numbers (S-LLM-CN-XGB in our benchmark). CARTE is is tailored to both strings and numbers.
An architecture that boosts performance
By using an architecture that models table entries not as a $i^{\text{th}}$ feature in a data matrix but as a function of its context (column names and neighboring entries) as well open-vocabulary embedding of strings, CARTE enables consistent representations of very different tables. This opens the door to pre-training across background tables, and fine-tuning to downstream tasks without matched entities or schema. Results show that after pretraining on a large knowledge base, the resulting model brings marked benefits to downstream analytic tasks, consistently outperforming a broad range of baselines both for learning on a single table or transfer learning on tables with imperfect correspondences. It enables transfer from tables in the wild, a setting so far never studied.
Toward tabular foundation models
Pre-training has been key to the wide application of deep learning on images and text. We hope that the ideas behind CARTE will bring these benefits to tabular learning, leading to tabular foundation models. This will call for further improved architectures: optimizing for larger train-sizes; refining numerical representations with ideas of Gorishniy et al. (2022); leveraging more training information and expressive attention as in large language models; merging with the complementary meta-learning ideas of TabPFN (Hollmann et al., 2023), which shines on heavily-numerical tables.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here. The societal impact of a method depends on how it is used. We do note, however, that tabular data are central to fields such as healthcare which uses a lot codes and more or less normalized entities (e.g., ICD10 codes for diseases, medical informatics as a field has invested hugely on data integration). We thus hope that pretraining a tabular model on health data could provide value to this field, and in turn positive societal impact.
On another topic, we note that our model, CARTE, comes with additional computational cost compared to baselines. We do expect that further research and engineering will bring these costs down. However, our work opens the door to pre-trained models for tabular data, one day maybe foundation models. These models have led to a race for ever-increasing size, which comes with dire consequences in terms of energy and financial cost, carrying over to ecology and concentration of power.
Acknowledgments
The authors acknowledge the support in part by the French Agence Nationale de la Recherche under Grant ANR-20-CHIA-0026 (LearnI).
We also would like to thank the ICML reviewers, who challenged us in a good way, leading to improved empirical study of CARTE, and hence a better manuscript.